# Does Generative AI speak Nigerian-Pidgin?: Issues about Representativeness and Bias for Multilingualism in LLMs

### v2

@@ 5.2 Machine translation evaluation
While statistical analysis already provided the difference between WAPE and Naija, evaluation on a task provides a different perspective i.e., would a model trained on WAPE, perform well on Naija, and vice versa? Also, what is the transfer performance of a model trained on one creole to another especially for text generation tasks?
Figure 2 shows our results on MT. The transfer performance depends on both the similarity of the domains (e.g. religious vs. news) and similarity of the creole style of writing. For example, religious datasets such as Bible and JW300 generally transfers poorly to other domains. Similarly, we find Mafand and UD in the news and spoken conversation domains also do not transfer well to religious domains.
If we focus on both Mafand and UD datasets, we find that they often have a higher zero-shot transfer result to WAPE (BBC and Wiki) than to Naija (Wiki). In the English-Pidgin direction, Mafand achieved between $76-80$ ChrF++ on WAPE while achieving only $52$ ChrF++ on Naija (PCM). Surprisingly, we find Mafand transfering better to WAPE than to its own test set showing the simplicity of generating WAPE pidgin compared to Naija. We have similar observation when transfering from UD. Our evaluation in the Pidgin-English also confirm this hypothesis that translating from WAPE to English is easier for MT systems than Naija.
@@ 5.3 LLM performance on Warri Benchmark
Here, we focus on finding out which pidgin is represented in the current LLMs, and whether they support several pidgin variants which are accessible for different communities of Naija speakers. We evaluated the performance of LLM in translating from and into WAPE and Naija.
@@
Our result (Table 4) of prompting GPT-4o, LLaMa 3.1 8B and LLaMa 3.1 70B shows that providing one or five examples is effective in providing more boost in performance for generating Pidgin sentences. [^10] The exact 1-shot and 5-shots examples are provided in the Appendix D For GPT-4o, the performance improved over zero-shot result by $+4.9$ ChrF++ when the LLM is prompted with one example translation of WAPE, and $+8.3$ ChrF++ when prompted with five examples, on the multi-way test set. However, the boost in performance is very small when Naija (Wiki) examples are provided. The boost is only $+2.6$ and $+3.0$ when one example and five examples are provided during the prompting of GPT-4o. This clearly shows that GPT-4o is more biased toward the WAPE than Naija and it is difficult to teach LLM with few examples. We believe one of the reason for this performance difference is because the WAPE BBC is the largest unlabelled data available on the web (Ogueji et al., 2021). Other sources that are more representative of Naija, are often in smaller quantity like Wikipedia. We observe similar trend for the LlaMa models where LLaMa 3.1 70B attained up to $80.3$ ChrF++ with 5-shots ($-2.7$ points when compared to GPT-4o), while LLaMa 3.1 8B achieved $76.4$.
@@
We provide a qualitative example in Table 5, where we show that with one or five examples, the LLaMa 3.1 70B LLM slightly changes its writing style to be more similar to Naija but sometimes the model combine the vocabulary of WAPE and Naija. For example, in the 5-shot translation of “…was expected of them to do in their different areas”, LLaMa 3.1 70B translated it to be “suppose do for dia different areas” which is more similar to the WAPE translation of the same sentence, while in Naija the words of “suppose”, “different” and “area” are spelt differently: “sopos du for dia difren aria”. On the other hand, GPT-4o almost produced an accurate translation in Naija except the use of “suppose” rather than “sopos” and “becos” (a WAPE word) instead of “bikos”. This implies that the model is able to learn in-context, however, it is still biased to WAPE without few-shot examples. With more examples, we may be able to teach the model Naija with supervised fine-tuning of instruction data containing Naija-English parallel sentences. Qualitative examples for WAPE, shows that the LLMs are able to correctly generate sentences in zero-shot setting without additionally few shot examples which confirms our hypothesis that the LLMs are biased to WAPE.
6 Qualitative interviews with Naija Wikipedia writers
@@
To validate our study, we interviewed two native speakers of Naija who contribute to the writing and editing of Naija Wikipedia articles. Some Wikipedian contributors have a public profile with links to their email addresses and social media accounts like Twitter or Linkedln. We sent emails to two Naija contributors with public profile and conducted a one hour interview with each of them.  [^11] We provided honorarium of \$$\$11$ to each interviewee.
@@
Through these interviews, we found that Naija Wikipedia writers are not linguists or language experts but they are just volunteers without a formal linguistic training. They have a passion for their Naija and make effort toward establishing a writing system which is very similar to the way they passive people speaking it in their community or in Nigeria, rather than trying to reach a wider audience like BBC. (2) They made a lot of effort to create standardized way of writing during the Naija Wikipedia incubator programme, otherwise, Wikipedia would not have included it as a separate language. These volunteer writers were part of the Wikipedia incubator program from the start, and they are part of the editors team of Naija Wikipedia. This team mentors new contributors about how to write the Naija reflecting the patterns how Naija is spoken (sometimes with a few adjustments to make it readable since original spoken Naija form could be different than English (e.g. “moto” instead of “car”)).
@@
The writers also mentioned that they read the available literature (e.g., Ofulue and Esizimetor (2010)) about Naija and follow the recommended rules by Naija linguists  Balogun (2013b); Aghoghovwia et al. (2010) before starting to contribute to Naija Wikipedia.
@@
In terms of content of the Naija articles, the writers focus on biographies of notable people (e.g., musicians and actors) in Nigeria. They are not allowed to contribute to sensitive topics like health except it is a direct translation from an high-resource language like English. The target audience for their articles is Naija speakers in Nigeria rather than the whole West Africa (targeted by BBC). To achieve this, they prefer words that come from local Nigerian languages (e.g., Hausa, Igbo and Yoruba), which many Nigerians are familiar with rather than words commonly understandable across West Africa.
@@ 7 Conclusion
Different versions of Pidgins mixed with English and local languages are used in West Africa but not all of them have standardized writing systems. Since 2017, BBC internet broadcasts in WAPE targets West African countries with the goal of reaching a wider audience in a standardized writing style.
Like many countries in West Africa, Nigeria is a multilingual country with both richness and challenges that come along with the diversity in the linguistic landscape. Although the official language is English, Naija is a lingua franca that brings speakers of different Nigerian languages together regardless of their linguistic, social or educational backgrounds. Since 2022, it is also a written language on Wikipedia with the efforts of volunteers.
Although both pidgin varieties are used in Nigeria, we prove that WAPE and Naija are very different from each other linguistically and current Generative AI models are built upon WAPE only. This is probably due to more availability of data on the Internet for the WAPE pidgin rather than Naija.
It is well-known that the lack of data on low resource languages is a key challenge for current AI systems which operate mostly on and for low resource languages. In our paper, we prove that the situation is much more challenging for linguistically rich areas (e.g., West Africa) where the pidgin varieties with the most data on the Internet gets represented on AI systems and the others are not visible. This could lead to bias towards favoring language preferences of certain speakers/users instead of being more inclusive toward the users/speakers of other pidgins.
@@ 8 Limitation
There are few limitations of our work (1) Our evaluation dataset is small, although we argue that 500 may be good enough as a test set for MT, however, we only have maximum of 5 sentences we could use for few-shot learning or in-context learning. Moreover, with additional sentences (e.g. 2.5K-5K parallel sentences as recommended in (Adelani et al., 2022)), we may be able to adapt M2M-100 model to produce better generation of the Wikipedia genre. (2) Our analysis is limited to one task which is machine translation, we hope to extend this analysis to other tasks in the future. (3) our analysis only cover Naija spoken in Nigeria, we hope to extend our analysis to other English-based creole in West Africa like Ghananian Pidgin, Cameronian Pidgin, and Krio in the future.

### v4

@@ 5.2 Machine translation evaluation
While statistical analysis already proves the linguistic difference between WAPE and Naija, evaluation on a task provides additional perspectives (i.e., Would a model trained on WAPE, perform well on Naija, and vice versa? What is the transfer performance of a model trained on one pidgin to another for text generation tasks?).
Figure 2 shows our results on MT. The transfer performance depends on both the similarity of the domains (e.g. religious vs. news) and similarity of the pidgin style of writing. For example, religious datasets (e.g., Bible and JW300) generally transfer poorly to other domains. Similarly, Mafand and UD in the news and spoken conversation domains also do not transfer well to religious domains.
Both Mafand and UD datasets often have a higher zero-shot transfer result to WAPE (BBC and Wiki) than to Naija (Wiki). In the English-Pidgin direction, Mafand achieved between $76-80$ ChrF++ on WAPE while achieving only $52$ ChrF++ on Naija (PCM). Surprisingly, we find Mafand transfering better to WAPE than to its own test set showing the simplicity of generating WAPE compared to Naija. We have a similar observation when transfering from UD. Our evaluation on the Pidgin-English also confirms this hypothesis that translating from WAPE to English is easier for MT systems than Naija.
@@ 5.3 LLM performance on Warri Benchmark
In this section, we focus on finding out which pidgin is represented in the current LLMs, and whether they support several pidgin variants which are accessible for different communities of Naija speakers. We evaluated the performance of LLM in translating from and into WAPE and Naija.
@@
Our result (Table 4) of prompting GPT-4o, LLaMa 3.1 8B and LLaMa 3.1 70B shows that providing one or five examples is effective for extra performance boost to generate generating Pidgin sentences. [^9] The exact 1-shot and 5-shots examples are provided in the Appendix D For GPT-4o, the performance improved over zero-shot result by $+4.9$ ChrF++ when the LLM is prompted with one example translation of WAPE, and $+8.3$ ChrF++ when prompted with five examples, on the multi-way test set. However, the boost in performance is very small when Naija (Wiki) examples are provided. It is only $+2.6$ and $+3.0$ when one example and five examples are provided during the prompting of GPT-4o. This shows that GPT-4o is more biased toward the WAPE than Naija and it is difficult to teach the LLM with few examples. The reason for this performance difference is because the WAPE (BBC) is the largest unlabelled data available on the web (Ogueji et al., 2021). Other sources that are more representative of Naija, are often in smaller quantity (e.g., Wikipedia). We observe a similar trend for the LlaMa models where LLaMa 3.1 70B attained up to $80.3$ ChrF++ with 5-shots ($-2.7$ points when compared to GPT-4o), while LLaMa 3.1 8B achieved $76.4$.
@@
We provide a qualitative example in Table 5, where we show that with one or five examples, the LLaMa 3.1 70B LLM slightly changes its writing style to be more similar to Naija but sometimes the model combine the vocabulary of WAPE and Naija which leads to misunderstandings. For example, in the 5-shot translation of “…was expected of them to do in their different areas”, LLaMa 3.1 70B translated it to be “suppose do for dia different areas” which is more similar to the WAPE translation of the same sentence. However, in Naija the words like “suppose”, “different” and “area” are spelled differently (e.g., “sopos du for dia difren aria”). On the other hand, GPT-4o produced an (almost) accurate translation into Naija except the use of “suppose” rather than “sopos” and “becos” (a WAPE word) instead of “bikos”. This implies that the model is able to learn in-context. However, it is still biased towards WAPE without few-shot examples. With more examples, we may be able to teach the model Naija with supervised fine-tuning of the instruction data containing Naija-English parallel sentences. Qualitative examples for WAPE show that the LLMs are able to generate sentences correctly in zero-shot setting without additionally few shot examples which confirms our hypothesis that the LLMs are biased towards WAPE.
6 Qualitative interviews with Naija Wikipedia contributors
@@
To validate our study, we interviewed two native speakers of Naija who contribute to the writing and editing of Naija Wikipedia articles. Some Wikipedian contributors have online public profiles with links to their email addresses and social media accounts (e.g, Twitter or Linkedln). We sent emails to two Naija contributors with the online public profiles and conducted interviews ($\sim 1$ hour each) with each of them.  [^10] We provided honorarium of \$$\$11$ to each interviewee.
@@
Our first observation is that the Naija Wikipedia contributors are not linguists or language experts but they are volunteers without a formal linguistic training. They have a passion for Naija and make an effort toward establishing a writing system which is very similar to the way it is spoken in their community (within Nigeria), rather than targeting a wider West African audience like BBC. They make efforts to create a standardized way of writing during the Naija Wikipedia incubator program. Through these efforts, Naija is included as a separate language on Wikipedia. The volunteers were part of the Wikipedia incubator program from the start, and they are part of the editors team of Naija Wikipedia. This team mentors new contributors about how to write the Naija reflecting the patterns how Naija is spoken (sometimes with a few adjustments to make it readable since original spoken Naija form could be different than English (e.g. “moto” instead of “car”)).
@@
The volunteers also mentioned that they consult the available literature (e.g., Ofulue and Esizimetor (2010)) about Naija and follow the recommended rules by Naija linguists  Balogun (2013b); Aghoghovwia et al. (2010) before starting to contribute to Naija Wikipedia.
@@
In terms of content of the Naija entries on Wikipedia, the contributors focus on the biographies of notable people (e.g., musicians and actors) in Nigeria. They are not allowed to contribute to sensitive topics (e.g., health) except when it is a direct translation from an high-resource language (e.g., English). To achieve this, they prefer words that come from local Nigerian languages (e.g., Hausa, Igbo and Yoruba), which many Nigerians are familiar with rather than words that are commonly understandable across West Africa. In general, the interviews with the Naija Wikipedia contributors confirm our results that they follow some convention distinguishing them from the WAPE writing convention.
@@ 7 Conclusion
Different versions of Pidgins mixed with English and local languages are used in West Africa but not all of them have standardized writing systems. Since 2017, BBC broadcasts (on Internet) in WAPE target the West African countries with the goal of reaching a wider audience across countries in a standardized writing style.
Nigeria is a multilingual country with both richness and challenges that come along with the linguistic diversity. Although the official language is English, Naija is a lingua franca that brings speakers of different Nigerian languages together regardless of their linguistic, social or educational backgrounds. Since 2022, it is also a written language on Wikipedia.
Although both pidgin varieties are used in Nigeria, we prove that WAPE and Naija are different from each other linguistically and current Generative AI models are built upon WAPE only. This is probably due to more availability of data on the Internet for the WAPE rather than Naija.
Lack of data on low resource languages is a key challenge for current AI systems. In our paper, we show that the situation is much more challenging for linguistically rich areas (e.g., West Africa). More specifically, pidgin varieties with the most data on the Internet gets represented on AI systems and the others may not be visible. This could potentially lead to a bias towards favoring language preferences of certain speakers/users instead of being more inclusive toward the users/speakers of other pidgins. Although our analysis focuses on Naija spoken in Nigeria, we hope to extend our analysis to other English-based pidgins in West Africa (e.g., Ghananian Pidgin, Cameronian Pidgin, and Krio in the future) as well.
@@ 8 Limitation
There are few limitations of our work (1) Our evaluation dataset is small, although we argue that 500 may be good enough as a test set for MT. However, we only have a maximum of 5 sentences we could use for the few-shot learning or in-context learning. Moreover, with additional sentences (e.g. 2.5K-5K parallel sentences as recommended in (Adelani et al., 2022)), we may be able to adapt M2M-100 model to produce better generation of the Wikipedia genre. (2) Our analysis is limited to one task which is machine translation, we hope to extend this analysis to other tasks in the future as well.
Acknowledgment
We are grateful to Sergiu Nisioi and John Pavlopoulos for their time and feedback on an early draft of the paper. We thank OpenAI for granting the API credits (through their Researcher Access API program to Masakhane) enabling the evaluation of GPT-4 LLMs. We thank the volunteer Wikipedia contributors for Naija for their insights and explanations, anonymous reviewers, area chairs and senior area chairs of NAACL25 for their feedback and recommendations on our paper.